Subject: APO 3.5m users committee minutes, 3/10/14

From: Suzanne Hawley

Submitted: Sat, 15 Mar 2014 15:27:24 -0700 (PDT)

Message number: 1328 (previous: 1327, next: 1329 up: Index)

 		APO 3.5-m Users Committee Phonecon, 3/10/14

Attending: Klaene, Ketzeback, Huehnerhoff, Stocke, Kuzio de Naray, Kaiser, 
Holtzman, Bowen, Indebetouw, Anderson, Hawley, Walterbos, Hughes Clark, 
Skrutskie, O'Connell, Kwitter

Absent: McMillan, Gillespie, Harper, Wisniewski 

User feedback and comments from institutional representatives:

o Washington - nothing to report.

o Chicago - no report.

o New Mexico State - nothing to report.

o Princeton - nothing to report.

o Colorado - nothing to report.

o Virginia - nothing to report.

o Johns Hopkins - nothing to report.

o Georgia State - nothing to report.

o Seattle University - Joanne said that during her last observing 
run there was an issue with the NA2 guider.

o NAPG - Karen said that her first remote observing night in January using 
DIS went well.

o University of Oklahoma - no report.


Discussion of telescope/instruments report:

Mark gave a brief summation of the telescope and instrument report.  He 
especially noted that during the upcoming months of March and April 
housing is very tight on the mountain.

Mark Klaene's detailed report for the reporting period follows:


 	      3.5-m Telescope and Instruments Highlights
 	                02/06/14 through 03/05/14

0) Overview

Generally a productive period with variable cloudiness. Snowfall amount 
has improved but is still only 25% of normal. March and April housing is 
way over subscribed doe to several classes, training sessions and SDSS IV 
activity. We are doing everything we can to keep students on site.

1) Telescope

Telescope has been working very well.  Progress continues to be made on 
the guider rewrite and some observers are using it.  We will continue to 
make progress on this over the next few months.

2) Instruments

DIS has been operating normally.

Agile is operational with no problems.

TripleSpec had no operational problems this period.

SPIcam variable dark current has been isolated to the current ion pump 
type and controller. Currently we see a spike that may affect dark current 
about 1/day and if it occurs during observations we will let the observer 
know to check their data.  Generally the effect is small. With warming 
weather coming we may be able to make progress on a long term solution 

Echelle is operational with no problems.

NIC-FPS continues to operate in shared-risk observing mode due to 
occasional image corruption issue.  NFS less than 9 is operational.

GIFS is operational. The small LN2 hold time continues to plaque us. The 
TUI interface is undergoing testing.

3) Summer shutdown planning is in work


o New users page:

The new users page is now linked directly from the APO home page, as the
first link under the 3.5m telescope section.  We have incorporated 
comments and feedback from the last meeting and from other users
who contacted us.  Suzanne thanked Alaina Bradley for her work on getting
this page up and running.  We welcome additional comments, especially
from new users.


o Q2 schedule:

The Q2 schedule has just been published and Suzanne pointed out a few 
items of note.  First, the GIFS on-site training run is scheduled for May 
27-28.  Chicago and Colorado users have expressed interest, and anyone 
else who would like to learn how to use GIFS, including hands-on 
experience with data acquisition in TUI and instruction on how to reduce 
the data cubes, is encouraged to contact Carol Grady and Mike McElwain by 
email (gifs-apo at and to attend the training run. 
More information on GIFS can be found at:

Suzanne also pointed out that the reservation policy for accommodations on 
the mountain has changed.  Users are now asked to put in their requests at 
least a month in advance if not more.  All users who are planning onsite 
visits in Q2 should be making their accommodation plans now.

There is a new monitoring program on the telescope this quarter for an 
hour a night for 45 nights.  This will impact the calibrations for first 
half users. John Stocke suggested that users should be in touch with each 
other about the possibility of sharing calibration data.

There is a program from a WIYN user that has been scheduled during Q2 in 
trade for time that APO users have received on the WIYN telescope in their 
semester 2014A.  WIYN users are being treated the same as those from ARC 
institutions; they must come to the telescope for 3 night training run and 
then can observe remotely.  Also with respect to WIYN time, Colorado and 
JHU users responded during Q2 scheduling with proposals for the WIYN 2014B 
semester.  Anyone else who wants time on WIYN during 2014B should send 
Suzanne a proposal on the APO cover page form.  She will be talking with 
WIYN management about time tradesfor 2014B and our quarters Q3 and Q4 in 
early April, so please send any proposals by April 1.  Joanne mentioned 
that she is observing on WIYN with ODI soon and will report back to the 
committee on her experience.  WIYN also requires that first time users 
travel to the site, and after that remote operations are possible.


o ARCTIC update:

Joe mentioned that he is working with STA to finalize the quote for the 
new CCD. We still need to come to a decision on the overcoating however 
that can take place after the order. He hopes to have the CCD in hand by 
September 1. There were a few questions regarding the coating and whether 
it would be blue sensitive and the response blue of 400nm. Joe responded 
that we are working with STA to develop a custom coating to give us the 
best performance for the widest general use. Any users who would like to 
be involved in the final selection of coating please contact either Mark 
or Suzanne.

Joe also mentioned that lifetime and light-tight lab testing on the 
prototype shutter is progressing. The optics design is moving forward and 
the goal is to complete work by May 1st.


o 20th anniversary celebration May 12-13:

Invitation lists were provided by the users committee members, and Rene 
said invitations will be mailed this week. Monday evening the 12th of May 
there will be a reception and then telescope viewing followed by APOLLO. 
Tuesday May 13 will include history and science talks with dinner at the 
Lodge following. In some cases, in order to save on postage, bundles of 
invitations will be coming to the users committee members addressed to 
others at their institutions. Be on the look out for these.  RSVPs are due 
by April 5.



[open from previous meeting]:

 	===> ACTION:  If users feel that the guider matching scripts are 
important, let Bill and Mark know.  Also, if there is someone willing to 
volunteer to help fix the scripts, likewise let Bill and Mark know.

Jack Dembicky has volunteered. He is not familiar with idl so it may take 
some time for this to get done. As a result we wanted to be clear on the 
priority to get this done and were looking for feedback from the users. 
Given we did not receive strong feedback we will give this a moderate 
priority.  This item is now closed.

 	===> ACTION:  Send expressions of interest in using ARCSAT in Q2 
to Suzanne

So far Suzanne has received one proposal through normal scheduling of 
3.5-m time. Suzanne will be sending out a call for proposals in April when 
the telescope has reached a state ready for shared-risk (no-cost) 
observing during Q2.  Testing of the remote interface is in process now, 
and science data were successfully obtained in early March with Joe's 
help. Karen Kwitter wanted to know about sending custom filters to the 
site.  She was told to contact Bill Ketzeback. It was noted by Joe that 
the filters should be verified to be less than 10 mm thick. Karen also 
wanted to do some testing in the f/8 beam. Suzanne said that should be 
fine.  Mark asked how often proposals would be accepted for ARCSAT. 
Suzanne replied that the current plan is to do scheduling by the quarter.

         ===> ACTION:  Send expressions of interest for WIYN time trades to 

See above discussion under Q2 schedule. Send proposals on APO cover page 
to Suzanne by April 1.  This action is open until April 1.

[new from this meeting]:



Next meeting:  The next Users Committee phonecon will be on Monday 14 
April at 8:30 AM Pacific Time.  The agenda and other materials will be 
sent to the committee members during the preceding week.

APO APO APO APO APO  Apache Point Observatory 3.5m  APO APO APO
APO  This is message 1328 in the apo35-general archive. You can find
APO  the archive on
APO  To join/leave the list, send mail to
APO  To post a message, mail it to